OpenStudy (anonymous):

Find the area of the rectangle. LENGTH: x+3 WIDTH: 2x+1 a. 2x2 + 6x + 3 b. 8x + 3 c. 2x2 + 7x + 3 d. 9x

OpenStudy (anonymous):

\[A = l \times w\] A=Area of the rectangle l=length w=width

OpenStudy (anonymous):

area is l*w = (x+3) (2x+1) then you just multiply the two: 2x^2+7x+3

OpenStudy (anonymous):

c

OpenStudy (anonymous):

therefore, in this case, \[A = (x+3)(2x+1)\] After, expand the brackets and get your answer.....

OpenStudy (anonymous):

okay then how about this one: If the length of a rectangular picture is (2x + 3) inches and the width is (x - 4) inches, which expression represents the area of the picture? a. (2x^2 + 5x - 12)in.^2 b. (2x^2 + 3x - 12)in.^2 c. (2x^2 - 8x - 12)in.^2 d. (2x^2 - 5x - 12)in.^2

OpenStudy (anonymous):

c?

OpenStudy (anonymous):

For your first question answer is option C And for your second question its option D

OpenStudy (anonymous):

(2x+3)(x-4)

OpenStudy (anonymous):

D